REANALYSIS OF THE NEAR INFRARED ELECTRONIC TRANSITIONS OF NiCl

\underline{LEAH C. O’BRIEN}, TAYLOR DAHMS, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; JAMES J. O’BRIEN, Department of Chemistry, University of Missouri, St Louis, MO 63121-4499.

Three near infrared electronic transitions of NiCl have been reanalyzed. Hougen’s recent (2011) theoretical investigation of low-lying states of metal halides with a d^{9} electronic configuration indicated that the signs of the experimental spinrotation and lambda-doubling signs for the low-lying Ω=1/2 states of NiCl were not correct. This problem is corrected by assuming an excited state with ^{2}Σ^{−} symmetry instead of ^{2}Σ^{+} symmetry. Line positions for the [12.3]^{2}Σ^{−} - X^{2}Π_{3/2}, [12.3]^{2}Σ^{−} - X^{2}Π_{1/2}, and [12.3]^{2}Σ^{−} - B^{2}Σ^{+} electronic transitions have been refitted, and the new spin-rotation and lambda-doubling signs agree with the Hougen theory. The results of the analysis will be presented.